OpenStudy (anonymous):

f(x) = 2x - 5 ------- 4x + 2 g(x) = 2x^2 + 7 Find: f(g(x)) (-- means its a fraction)

OpenStudy (anonymous):

\[f(g(x)) = \frac{ 2g(x)-5 }{ 4g(x)+2 }=\frac{ 4x^{2}+9 }{ 8x^{2}+16 }\]